OpenStudy (anonymous):

how do you graph an equation like this: y-1=-2/3(x-6)?

OpenStudy (anonymous):

Add 1 to both sides y = 1/3(x-6) Multiply both sides by 3 3y = x-6 Add 6 to both sides 3y + 6 = x Let x = 0 then LHS makes 3y = -6 So when x - 0 y = -2 Plot this coordinate (0, -2) Now go back to the equation 3y + 6 = x and let y = 0 making x =6 Now plot that coordinate (6, 0) Join a line through the two coordinates voila!